# An evaluation of LLMs and Google Translate for translation of selected Indian languages via sentiment and semantic analyses

### v2

@@ Abstract
Large Language models (LLMs) have been prominent for language translation, including low-resource languages. There has been limited study about the assessment of the quality of translations generated by LLMs, including Gemini, GPT and Google Translate. In this study, we address this limitation by using semantic and sentiment analysis of selected LLMs for Indian languages, including Sanskrit, Telugu and Hindi. We select prominent texts that have been well translated by experts and use LLMs to generate their translations to English, and then we provide a comparison with selected expert (human) translations. Our findings suggest that while LLMs have made significant progress in translation accuracy, challenges remain in preserving sentiment and semantic integrity, especially in figurative and philosophical contexts. The sentiment analysis revealed that GPT-4o and GPT-3.5 are better at preserving the sentiments for the Bhagavad Gita (Sanskrit-English) translations when compared to Google Translate. We observed a similar trend for the case of Tamas (Hindi-English) and Maha P (Telugu-English) translations. GPT-4o performs similarly to GPT-3.5 in the translation in terms of sentiments for the three languages. We found that LLMs are generally better at translation for capturing sentiments when compared to Google Translate.
journal: .
@@
[…]
[inst3]Centre for Artificial Intelligence and Innovation, Pingla Institute, Sydney, Australia
1 Introduction
@@
Recent developments in utilising deep learning models [1] in natural language processing (NLP) [2] had a significant impact on the field of machine translation (language translation) [3, 4, 5]. This led to the development of sophisticated models [6] capable of understanding and translating text, such as Google Translate [7].
@@
Machine translation systems [8, 9] refer to translation software systems that have achieved remarkable success in recent years due to multilingual pre-trained language models [10]. Large Language models (LLMs) [11, 12, 13] are pre-trained deep learning models that have emerged as a promising tool for translation systems and other NLP tasks such as multilingual machine translation [14], text summarizations [15] and language modelling [16]. However, the translation quality of these models, especially for low-resource languages [17], often falls short of that achieved by models trained on higher-resource languages. Low resource languages refer to the less studied, resource-scarce, less computerised, less commonly taught or low-density languages [18, 19, 20]. The scarcity of extensive monolingual or parallel corpora poses significant challenges to the development of accurate and culturally sensitive machine translation systems [21, 22].
@@
Natural Language Generation (NGL) [23] enables computational devices to understand and write text that is indistinguishable from the work of humans across a vast array of languages. Gemini is a multimodal LLM developed by Google DeepMind and released in November 2023 which is a predecessor of PaLM (Pathways Language Model) [24]. Gemini is based on variants of Generalised Pre-trained Transformer (GPT) models [25, 26, 27] which was originally released as Chat-GPT in November 2022 by OpenAI [28]. Chat-GPT enables a conversation with user prompts and replies, which can also be used for language translation [29]. Although prominent, GPT-based models are criticised for lack of accountability and transparency [30, 31], explainability [32] and expertise in a specific field, i.e. they have been called ’master of none’ [33]. Gemini and Chat-GPT have brought an entirely new era to machine translation [34, 35] which is being validated by comparing with expert (human) translators [36]. However, validating low-resource languages can be challenging due to constraints in the amount of data available [37]. The use of transfer learning[38], multilingual models[39], and hybrid systems [40] can improve the quality of translations.
@@
Low-resource languages such as Telugu, Hindi and Sanskrit are often overlooked in the realm of language technology including translation, speech recognition and NLP problems [41]. These languages are characterised by their under-representation in digital datasets and a scarcity of linguistic resources. Hence, they present unique challenges for researchers and developers aiming to create effective language processing systems [42]. The complexity of these languages, coupled with the limited availability of resources, requires innovative approaches to translation and language processing. Hendy et al. [29] reported that GPT models are better at translating high-resource models but are lacking when it comes to low-resource languages. Similarly, an approach by Sethi et al. [43] applied GPT models for Sanskrit to Hindi translation, highlighting the potential of GPT models to enhance translation quality despite the scarcity of resources. These studies underscore the importance of developing innovative methods to address the challenges posed by low-resource languages.
@@
A critical aspect of our study is the analysis of sentiment and semantic elements in the translated text. Sentiment analysis [44, 45] involves detecting the emotional tone expressed, typically via deep learning and LLMs. Chandra et al.[46] presented an evaluation of Google Translate for Sanskrit to English translation via sentiment and semantic analysis. They made a framework to evaluate the translation of Sanskrit by Google Translate by sentiment and semantic analysis. They analysed the translation of the Bhagavad Gita (Sanskrit to English), by Google Translate and compared it with expert (human) translations, which motivates our study.
@@
Semantic analysis [47, 48], on the other hand, focuses on the meaning of words and phrases, ensuring that the translated text accurately conveys the intended meaning [49]. By looking at how precise the translations are semantically and whether they remain faithful to the original meaning. Hence, we can compare LLM translations with expert (human) translations using similarity scores such as the BLEU (Bilingual Evaluation Understudy) score [50].
@@
In the past decades, there has been much focus on the analysis of the quality of machine translation, which is crucial for the advancement and practical implementation of translation applications. Gamon et al. [51] developed an automatic evaluation process using BLEU and compared it with human translations via n-grams as an indirect indicator of translation quality. Hendy et al. [29] used GPT models for machine translation tasks by focusing on the zero-shot translation capabilities of GPT models, which are capable of performing well on multiple language pairs without any pre-training procedures or domain-specific customisation process [52]. This breakdown reveals their intrinsic language acquisition and expression abilities and the effect of making use of various prompting mechanisms on the way GPT models are translated. Zhang et al. [53] reviewed different types of prompts that trigger the most authentic and fluent translation produced by these language models to estimate the capacity of GPT models for document-level translation. Yang et al. [54] examined the effects of GPT and the traditional neural machine translation models and reported their synergy provides much higher quality.
@@
In this study, we evaluate the potential of LLMs in translating selected low-resource Indian languages, using sentiment and semantic analysis. We use Google Translate, Gemini, and GPT to translate prominent texts that are prominent in the region and translated by experts to provide a comparison with expert (human) translations. In order to validate the effectiveness of the translations, we employ different methodologies including lexicon-based sentiment analysis and semantic analysis for comparison and evaluation of translation quality. We select Sanskrit, Telugu and Hindi texts for our study and provide translated versions for further analysis via open-source code and data. Our research includes low-resource languages that are spoken by millions of people in the hope of preserving the beauty of human expression and looking for new ways to connect people from different cultures of the world.
@@
2 Background
2.1 Sentiment and Semantic Analysis
@@
Sentiment analysis and semantic analysis are two critical areas in NLP that aim to understand and interpret human language. Sentiment analysis [60] also known as opinion mining, focuses on determining the sentiment or emotion expressed in a piece of text, such as whether the text conveys positive, negative, and neutral sentiments [61]. Sentiment and semantic analysis become mutually complementary in a manner that allows for a complete interpretation of opinions, sentiments, entities, relations and the meaning of the surrounding context in the given text data [62]. Implementing transfer learning and transformer-based language models has significantly improved the accuracy of implementing such systems [63]. Nonetheless, Misra et al. [64], introduced a large-scale and high-quality dataset and framework for deep learning models that can reliably learn sarcastic cues from the text in an expressive manner. In the case of our study, a combination of sentiment and semantic analysis would be useful as done in earlier studies [38, 65].
@@
Sentiment analysis can be done at three levels: sentence level, document level and aspect level. At the sentence level, the document consisting of paragraphs is broken into sentences and sentiment analysis is done for every sentence [66]. Document-level sentiment analysis consists of the sentiment extraction from the entire document, intended for gaining global sentiment so that there will be no local unnecessary patterns[67]. Differentiating the sharp distinction between the intricate internals of sentiments and the dependent words at this level demands greater insight. In the aspect level sentiment analysis, emphasis is placed on opinions concerning the features or attributes that are expressed in the text [68]. LSTM models with attention mechanism [69] addressed this difficulty by only distinguishing multiple aspects in a sentence, but eventually measuring the polarity of each aspect at the same time.
@@
The combination of sentiment analysis and semantic analysis techniques has become a powerful tool for mining comprehension of human language data at deeper levels [70]. The main obligations of sentiment analysis lie in the classification of overall perceptions and feelings, yet semantic analysis concerns itself with contextual meaning, linguistic relationships and interpretation of the lexemes [71]. The semantic analysis aims to derive meaning representations from text [72] which encompasses a diverse set of techniques to uncover the underlying semantics, including named entity recognition, coreference resolution, semantic role labelling, and relationship extraction [73]. Chandra et al. [65] used a combination of sentiment and semantic analysis to compare different translations of the Bhagavad Gita (Sanskrit - English) which was later compared with Google Translate [74]. These studies motivate their combination to be used for comparison of LLMs for translation of selected Indian languages.
2.2 BERT
@@
The BERT (Bidirectional Encoder Representations from Transformers) model [75] provides a significant advancement in NLP. It employs a masked language modelling approach [76] to analyse the whole text from both directions, providing more contextual understanding and providing better word embedding in NLP tasks such as sentiment analysis. The awareness of certain contexts is important for translation because the sense of a word may change upon emerging in different texts [77]. In the case of machine translation using BERT, the context apprehension of BERT allows the model to accept literary devices such as idioms, and allegory [76]. These usually confuse traditional machine translation systems that can’t grasp or understand the context. This facilitates better translations that not only mean the beginning but also understand the tone and nuance [78]. Hence, BERT-based models have gained immense attention in NLP applications, including sentiment analysis [79]. Prominent BERT-based implementations include advanced models (MPNet [80]), hate speech analysis (HateBERT [81]), and semantic analysis (Semantic-BERT[82]). Recently, Singh and Chandra released the Hindubhobic-BERT model [83] for sentiment and abuse analysis of Hindubhobic text based on X (Twitter). In our study, we will use BERT-based models for sentiment and semantic analysis as done in previous studies [65, 74]
3 Methodology
3.1 Data
@@
According to 2011 Indian Census, Hindi is spoken by more than 43 % of the Indian population as first and second languages  [^1] https://web.archive.org/web/20180627091302/http://www.censusindia.gov.in/2011Census/C-16_25062018_NEW.pdf. Taking into account 2024 population projection by Worldometer  [^2] https://www.worldometers.info/world-population/india-population/, we produce Table 1 that shows the percentage of Hindi, Telugu and Sanskrit speakers in 2011 and 2024. Although Hindi and Telugu feature large language populations, they would still be considered low-resource languages, due to the use of these languages in the education system, resources available in language translation, and analyses of translations. Next, we present the selected texts for the three languages for our study.
@@
Table 1: Projected population of selected languages in India (2011 vs 2024).
3.1.1 Sanskrit: Bhagavad Gita
@@
The Bhagavad Gita [84] [85, 86] is one of the foundational texts of Hinduism, encompassing practices, spirituality and philosophy. The text features eighteen chapters and is divided into several parts, each of them addressing various philosophical topics, such as duty (dharma), action (karma), spiritual way of life, and the nature of the self (Atman). The first chapter introduces the protagonists (Arjuna and Krishna) and the problem setting of a philosophical nature based on an event from the Mahabharata, i.e. Arjuna’s ethical dilemma to go or not to go to war for dharma (duty).
@@
The original language of the Bhagavad Gita is Sanskrit, featuring 701 verses (sloka) describing a conversation from the Mahabharata war in India. The Bhagavad Gita is a synthesis of various Hindu philosophies, including Vedanta, Yoga and Samkhya and discusses and synthesises the three dominant themes in Hinduism: enlightenment-based renunciation, dharma-based household life and devotion-based theism[84]. The Bhagavad Gita over the past two millenniums and has been translated numerous times into hundreds of languages with influence worldwide [87]. It has been studied with perspectives of leadership [88], psychology and psychotherapy [89, 86], and management [90]. As noted earlier, Shukla et al. [74] translated the Bhagavad Gita into English using Google Translate and compared different translations using sentiment analysis. Currently, Sanskrit is thriving in academia given that the majority of the Hindu texts were originally written in Sanskrit, but it struggling as a language of everyday use and communication as shown in Table 1.
3.1.2 Telugu: Maha Prasthanam
@@
Maha Prasthanam [91] [^3] https://greaterTelugu.org/sri-sri-mahaprasthanam-pdf-book/, is a Telugu-language anthology of poems by the noted literary writer Srirangam Srinivasarao which features 18 chapters. It is this anthology which is considered an epic and a magnum opus in modern Indian poetry that was published in 1950 and ever since it has completely transformed the Telugu drama. The book is a poem collection which was written between 1930 and 1940 and when it was published, a milestone was made in Telugu literature. The title (Maha Prasthanam) is an amalgamation of two words, prasthanam meaning ’the way’ and maha, the superlative for ’great’. This is an apt description of the depth of thematic and philosophical exploration in the poem. It has been translated into English by K. S. Sastry in 1979 [91] which made the work more accessible and thus makes it possible to analyse it to bridge the language gap between Telugu and English.
3.1.3 Hindi: Tamas

### v3

@@ Abstract
Large Language models (LLMs) have been prominent for language translation, including low-resource languages. There has been limited study on the assessment of the quality of translations generated by LLMs, including Gemini, GPT, and Google Translate. This study addresses this limitation by using semantic and sentiment analysis of selected LLMs for Indian languages, including Sanskrit, Telugu and Hindi. We select prominent texts (Bhagavad Gita, Tamas and Maha Prasthanam ) that have been well translated by experts and use LLMs to generate their translations into English, and provide a comparison with selected expert (human) translations. Our investigation revealed that while LLMs have made significant progress in translation accuracy, challenges remain in preserving sentiment and semantic integrity, especially in metaphorical and philosophical contexts for texts such as the Bhagavad Gita. The sentiment analysis revealed that GPT models are better at preserving the sentiment polarity for the given texts when compared to human (expert) translation. The results revealed that GPT models are generally better at maintaining the sentiment and semantics when compared to Google Translate. This study could help in the development of accurate and culturally sensitive translation systems for large language models.
Index Terms:
LLMs, sentiment analysis, machine translation, and semantic analysis.
@@
=-15pt
I Introduction
@@
Recent developments in the use of deep learning models [1] in natural language processing (NLP) [2] had a significant impact on the field of machine translation (language translation) [3, 4, 5]. This led to the development of sophisticated models [6] capable of understanding and translating text, such as Google Translate [7].
@@
Machine translation systems [8, 9] refer to translation software systems that have achieved remarkable success in recent years due to multilingual pre-trained language models [10]. Large Language models (LLMs) [11, 12, 13] are pre-trained deep learning models that have emerged as a promising tool for translation systems and other NLP tasks such as multilingual machine translation [14], text summarisation [15] and language modelling [16]. However, the translation quality of these models, especially for low-resource languages [17], often falls short of that achieved by models trained on higher-resource languages. Low resource languages refer to the less studied, resource-scarce, less computerised, less commonly taught or low-density languages [18, 19, 20]. The scarcity of extensive monolingual or parallel corpora poses significant challenges to the development of accurate and culturally sensitive machine translation systems [21, 22].
@@
Natural Language Generation (NLG) [23] enables computational devices to understand and write text that is indistinguishable from the work of humans across a vast array of languages. Gemini is a multimodal LLM developed by Google DeepMind and released in November 2023, which is a predecessor of PaLM (Pathways Language Model) [24]. Gemini is based on variants of Generalised Pre-trained Transformer (GPT) models [25, 26, 27] that were originally released as Chat-GPT in November 2022 by OpenAI [28]. Chat-GPT enables a conversation with user prompts and replies, which can also be used for language translation [29]. Although prominent, GPT-based models are criticised for a lack of accountability and transparency [30, 31], explainability [32], and expertise in a specific field, i.e. they have been called ’master of none’ [33]. Gemini and Chat-GPT have brought an entirely new era to machine translation [34, 35], which is being validated by comparing with expert (human) translators [36]. However, validating low-resource languages can be challenging due to constraints in the amount of data available [37]. The use of transfer learning[38], multilingual models[39], and hybrid systems [40] can improve the quality of translations.
@@
Low-resource languages such as Telugu, Hindi and Sanskrit are often overlooked in the realm of language technology, including translation, speech recognition and NLP problems [41]. These languages are characterised by their under-representation in digital datasets and a scarcity of linguistic resources. Hence, they present unique challenges for researchers and developers aiming to create effective language processing systems [42]. The complexity of these languages, coupled with the limited availability of resources, requires innovative approaches to translation and language processing. Hendy et al. [29] reported that GPT models are better at translating high-resource models but are lacking when it comes to low-resource languages. Similarly, an approach by Sethi et al. [43] applied GPT models for Sanskrit to Hindi translation, highlighting the potential of GPT models to enhance translation quality despite the scarcity of resources. These studies underscore the importance of developing innovative methods to address the challenges posed by low-resource languages.
@@
A critical aspect of our study is the analysis of sentiment and semantic elements in the translated text. Sentiment analysis [44, 45] involves detecting the emotional tone expressed, typically via deep learning and LLMs. Chandra et al.[46] presented an evaluation of Google Translate for Sanskrit to English translation via sentiment and semantic analysis. The study evaluated the translation quality of Sanskrit-English by Google Translate using sentiment and semantic analysis. Their study analysed the translation of the Bhagavad Gita by Google Translate and compared it with expert (human) translations, which motivates our study.
@@
Semantic analysis [47, 48], on the other hand, focuses on the meaning of words and phrases, ensuring that the translated text accurately conveys the intended meaning [49]. Semantic analysis can check if the translation remains faithful to the original text in terms of meaning. Hence, we can compare LLM translations with expert (human) translations using similarity scores such as the BLEU (Bilingual Evaluation Understudy) [50].
@@
In the past decades, there has been much focus on the analysis of the quality of machine translation, which is crucial for the advancement and practical implementation of translation applications. Gamon et al. [51] developed an automatic evaluation process using BLEU and compared it with human translations via n-grams as an indirect indicator of translation quality. Hendy et al. [29] used GPT models for machine translation tasks by focusing on the zero-shot translation capabilities of GPT models, which are capable of performing well on multiple language pairs without any pre-training procedures or domain-specific customisation process [52]. This breakdown reveals their intrinsic language acquisition and expression abilities and the effect of making use of various prompting mechanisms on the way GPT models are translated. Zhang et al. [53] reviewed different types of prompts that trigger the most authentic and fluent translation produced by these language models to estimate the capacity of GPT models for document-level translation. Yang et al. [54] examined the effects of GPT and the traditional neural machine translation models, and reported that their synergy provides much higher quality.
@@
In this study, we evaluate the potential of LLMs in translating selected low-resource Indian languages, using sentiment and semantic analysis. We use Google Translate, Gemini, and GPT-4o to translate prominent texts that are prominent in the region and translated by experts to provide a comparison with expert (human) translations. In order to validate the effectiveness of the translations, we employ different methodologies, including lexicon-based sentiment analysis and semantic analysis for comparison and evaluation of translation quality. We select Sanskrit, Telugu and Hindi texts for our study and provide translated versions for further analysis via open-source code and data. Our research includes low-resource languages that are spoken by millions of people in the hope of preserving the beauty of human expression and looking for new ways to connect people from different cultures around the world.
@@
II Background
II-A Sentiment and Semantic Analysis
@@
Sentiment analysis and semantic analysis are two critical areas in NLP that aim to understand and interpret human language. Sentiment analysis [60], also known as opinion mining, focuses on determining the sentiment or emotion expressed in a piece of text, such as whether the text conveys positive, negative, and neutral sentiments [61]. Sentiment and semantic analysis become mutually complementary in a manner that allows for a complete interpretation of opinions, sentiments, entities, relations and the meaning of the surrounding context in the given text data [62]. Implementing transfer learning and transformer-based language models has significantly improved the accuracy of implementing such systems [63]. Nonetheless, Misra et al. [64] introduced a large-scale and high-quality dataset and framework for deep learning models that can reliably learn sarcastic cues from the text in an expressive manner. In the case of our study, a combination of sentiment and semantic analysis would be useful, as done in earlier studies [38, 65].
@@
Sentiment analysis can be done at three levels: sentence level, document level and aspect level. At the sentence level, the document consisting of paragraphs is broken into sentences and sentiment analysis is done for every sentence [66]. Document-level sentiment analysis consists of the sentiment extraction from the entire document, intended for gaining global sentiment so that there will be no local unnecessary patterns[67]. Differentiating the sharp distinction between the intricate internals of sentiments and the dependent words at this level demands greater insight. In the aspect-level sentiment analysis, emphasis is placed on opinions concerning the features or attributes that are expressed in the text [68]. LSTM models with attention mechanism [69] addressed this difficulty by only distinguishing multiple aspects in a sentence, but eventually measuring the polarity of each aspect at the same time.
@@
The combination of sentiment analysis and semantic analysis techniques has become a powerful tool for mining comprehension of human language data at deeper levels [70]. The main obligations of sentiment analysis lie in the classification of overall perceptions and feelings, yet semantic analysis concerns itself with contextual meaning, linguistic relationships and interpretation of the lexemes [71]. The semantic analysis aims to derive meaning representations from text [72], which encompasses a diverse set of techniques to uncover the underlying semantics, including named entity recognition, coreference resolution, semantic role labelling, and relationship extraction [73]. Chandra et al. [65] used a combination of sentiment and semantic analysis to compare different translations of the Bhagavad Gita (Sanskrit-English), which was later compared with Google Translate [74]. These studies motivate their combination to be used for the comparison of LLMs for the translation of selected Indian languages.
II-B BERT
@@
The BERT (Bidirectional Encoder Representations from Transformers) model [75] provides a significant advancement in NLP. It employs a masked language modelling approach [76] to analyse the whole text from both directions, providing more contextual understanding and better word embedding in NLP tasks such as sentiment analysis. The awareness of certain contexts is important for translation because the sense of a word may change upon emerging in different texts [77]. In the case of machine translation using BERT, the context apprehension of BERT allows the model to accept literary devices such as idioms, and allegory [76]. These usually confuse traditional machine translation systems that can’t grasp or understand the context. This facilitates better translations that not only mean the beginning but also understand the tone and nuance [78]. Hence, BERT-based models have gained immense attention in NLP applications, including sentiment analysis [79]. Prominent BERT-based implementations include advanced models (MPNet [80]), hate speech analysis (HateBERT [81]), and semantic analysis (Semantic-BERT[82]). Recently, Singh and Chandra released the Hindubhobic-BERT model [83] for sentiment and abuse analysis of Hindubhobic text based on X (Twitter). In our study, we will use BERT-based models for sentiment and semantic analysis as done in previous studies [65, 74]
III Methodology
III-A Data
@@
According to 2011 Indian Census, Hindi is spoken by more than 43 % of the Indian population as first and second languages  [^1] https://web.archive.org/web/20180627091302/http://www.censusindia.gov.in/2011Census/C-16_25062018_NEW.pdf. Taking into account 2024 population projection by Worldometer  [^2] https://www.worldometers.info/world-population/india-population/, we produce Table I that shows the percentage of Hindi, Telugu and Sanskrit speakers in 2011 and 2024. Although Hindi and Telugu feature large language populations, they would still be considered low-resource languages, due to the use of these languages in the education system, resources available in language translation, and analyses of translations. We remove the stop words and undesired characters for bigrams and trigram analysis. In the case of translation by LLMs, we only remove the numbers featured in the respective texts. Next, we present the selected texts for the three languages for our study.
@@
TABLE I: Projected population of selected languages in India (2011 vs 2024).
III-A1 Sanskrit: Bhagavad Gita
@@
The Bhagavad Gita [84] [85, 86] is one of the foundational texts of Hinduism, encompassing practices, spirituality and philosophy. The text features eighteen chapters and is divided into several parts, each of which addresses various philosophical topics, such as duty (dharma), action (karma), spiritual way of life, and the nature of the self (Atman). The first chapter introduces the protagonists (Arjuna and Krishna) and the problem setting of a philosophical nature based on an event from the Mahabharata, i.e. Arjuna’s ethical dilemma to go or not to go to war for dharma (duty).
@@
The original language of the Bhagavad Gita is Sanskrit, featuring 701 verses (sloka) describing a conversation from the Mahabharata war in India. The Bhagavad Gita is a synthesis of various Hindu philosophies, including Vedanta, Yoga and Samkhya and discusses and synthesises the three dominant themes in Hinduism: enlightenment-based renunciation, dharma-based household life and devotion-based theism[84]. The Bhagavad Gita has been translated over the past two millennia and has been translated numerous times into hundreds of languages, with influence worldwide [87]. It has been studied with perspectives of leadership [88], psychology and psychotherapy [89, 86], and management [90]. As noted earlier, Shukla et al. [74] translated the Bhagavad Gita into English using Google Translate and compared different translations using sentiment analysis. Currently, Sanskrit is thriving in academia, given that the majority of the Hindu texts were originally written in Sanskrit, but it is struggling as a language of everyday use and communication, as shown in Table I.
III-A2 Telugu: Maha Prasthanam
@@
Maha Prasthanam [91] [^3] https://greaterTelugu.org/sri-sri-mahaprasthanam-pdf-book/ is a Telugu-language anthology of poems by the noted literary writer Srirangam Srinivasarao, which features 18 chapters. It is this anthology that is considered an epic and a magnum opus in modern Indian poetry, which was published in 1950, and ever since, it has completely transformed the Telugu drama. The book is a poem collection which was written between 1930 and 1940, and when it was published, a milestone was made in Telugu literature. The title (Maha Prasthanam) is an amalgamation of two words, prasthanam meaning ’the way’ and maha, the superlative for ’great’. This is an apt description of the depth of thematic and philosophical exploration in the poem. It has been translated into English by K. S. Sastry in 1979 [91], which made the work more accessible and thus makes it possible to analyse it to bridge the language gap between Telugu and English.
III-A3 Hindi: Tamas